OpenStudy (anonymous):

can some give me the steps to solve this problem 7e^4x+1=28

OpenStudy (anonymous):

First you want to subtract the 1 over and then divide by 7. So you will have e^4x=27/7 Then to get rid of the e, you natural log both sides. So you get 4x=ln(27/7) Then finally, x=(ln(27/7))/4

OpenStudy (anonymous):

okay, I was dividing both sides by 7 first

OpenStudy (anonymous):

That would work, you'd just have to make sure that you also divide that 1 by 7. It's usually just simpler to combine constants first.